# Cardinality-Preserving Attention Channels for Graph Transformers in Molecular Property Prediction

### v3

@@ Abstract
Drug discovery motivates accurate molecular property prediction when labeled data are limited and candidate spaces are vast. This article presents CardinalGraphFormer, a graph transformer that augments structured attention with a query-conditioned gated unnormalized aggregation channel to preserve dynamic cardinality signals, complemented by graph-specific structural biases; a locality prior via sparse masking provides scalability for larger graphs. For typical drug-like molecules ($K=3$ is near-global), masking acts mainly as a regularizer; for larger graphs it provides meaningful efficiency gains. Pretraining unifies contrastive alignment of augmented graph views and masked reconstruction of attributes. Evaluations on public benchmarks show consistent gains over baselines, isolated via controls for capacity, objectives, and size effects. Ablations confirm the cardinality channel’s contributions beyond simpler approximations, with efficiency benefits on large molecules. Code, artifacts, and protocols emphasize reproducibility.
@@ 1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [10, 2]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [17, 18, 5]. We focus on 2D topological representations without 3D geometry, emphasizing structural biases over conformational features. We adopt Graphormer-inspired structural biases (shortest-path distance (SPD) + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks as a locality prior for scalability on larger graphs. We treat this mask primarily as a locality prior and efficiency knob; the main methodological contribution is the cardinality-preserving attention (CPA) channel, which retains dynamic support-size signals complementary to static centrality embeddings.
@@
  • A graph transformer layer with query-conditioned cardinality-preserving attention aligned to structured supports.
  • Dual-objective self-supervised pretraining on $\sim$28M molecules.
  • Rigorous matched-protocol evaluation showing consistent, often significant gains over reproduced baselines, expanded to include Graphormer, D-MPNN, GIN, and contextualized against pretrained 2D models (GROVER, MolCLR, GEM). Code and artifacts will be released upon acceptance (reserved DOI: 10.5281/zenodo.18622116).
@@ 2 Related Work
Message passing remains a dominant paradigm for molecular graphs [4], and attention-based GNNs are widely used [12]. Graph transformers such as Graphormer inject structural biases (shortest-path distance, edge features, centrality) into global attention [17]. Sparse variants often restrict attention to local neighborhoods (e.g., $K$-hop masking) to reduce computation in practice. Cardinality-preserving attention (CPA) adds unnormalized channels to retain degree/cardinality signals [20]. Self-supervised pretraining on graphs includes contrastive methods (GraphCL) and masked modeling (GraphMAE) [18, 5]. CardinalGraphFormer unifies structured sparse attention with aligned cardinality preservation and dual-objective pretraining.
@@ 3.2 Structured Sparse Attention with Graphormer-Inspired Biases
The support set is $\mathcal{S}(i)=\{j\mid\mathrm{SPD}(i,j)\leq K\}$ (includes self, SPD(i,i)=0(i,i)=0). Main experiments use $K=3$. Attention is masked to $\mathcal{S}(i)$ (outside: -$\infty$). Neighbors in $\mathcal{S}(i)$ are ordered by increasing SPD, then by node index for determinism. Logits:
@@
We implement structured attention by precomputing neighbor index lists per node and using gather-attention-scatter operations over the support set. To remain compatible with optimized attention kernels (e.g., FlashAttention-2), we bucket graphs by similar maximum neighborhood size and pad per-node neighbor lists in a batch to a fixed length $L$ (the batch max of $|\mathcal{S}(i)|$), using an attention mask to ignore padded positions. Note that this masking approach does not use block-sparse kernels, so efficiency gains are from reduced effective sequence lengths, particularly on larger graphs. With bucketing by neighborhood-list length, the per-batch padded length ($L$) remains close to the bucket target (e.g., mean $L\approx 25$ on the full corpus), yielding ¡10% average padding.
@@ 3.3 Cardinality-Preserving Attention
The degree-normalized CPA ablation (Table 13) divides the unnormalized sum by $|\mathcal{S}(i)|$ before gating and underperforms, consistent with the hypothesis that raw magnitude variation—correlated with current support size—carries useful information beyond static degree bins or explicit normalization. Supporting this, the L2 norm of the gated unnormalized sum (averaged across heads and layers, post-LayerNorm) correlates positively with $|\mathcal{S}(i)|$ even after controlling for node degree (partial correlation $r\approx 0.52$), computed on a held-out 50k-graph pretraining validation subset, averaged across 5 seeds.
@@ 3.4 Self-Supervised Objectives
Contrastive views. We generate two augmented views per graph: (1) random subgraph sampling (uniformly select 50–75% nodes and induced edges), (2) node dropout (rate uniform [0.1,0.3] on surviving nodes), (3) edge dropout (rate uniform [0.1,0.3] on surviving edges). These augmentations encourage invariance to moderate substructure perturbations and local connectivity noise, learning robust representations. To address conceptual mismatch concerns, we now recompute SPD distances and degree bins per view, intersecting $\mathcal{S}(i)$ with surviving nodes. This increases computation but enhances semantic consistency; ablations (Table 14) confirm improved downstream performance (e.g., +0.005–0.01 on select metrics) over stable SPD (original design) and attribute-only augmentations. We also ablate chemistry-aware alternatives (attribute masking, valency-constrained edge perturbations), showing CPA gains persist (Table 1).

### v4

@@ Abstract
Molecular property prediction is crucial for drug discovery when labeled data are scarce. This work presents CardinalGraphFormer, a graph transformer augmented with a query-conditioned cardinality-preserving attention (CPA) channel that retains dynamic support-size signals complementary to static centrality embeddings. The approach combines structured sparse attention with Graphormer-inspired biases (shortest-path distance, centrality, direct-bond features) and unified dual-objective self-supervised pretraining (masked reconstruction and contrastive alignment of augmented views). Evaluation on 11 public benchmarks spanning MoleculeNet, OGB, and TDC ADMET demonstrates consistent improvements over protocol-matched baselines under matched pretraining, optimization, and hyperparameter tuning. Rigorous ablations confirm CPA’s contributions and rule out simple size shortcuts. Code and reproducibility artifacts are provided.
@@ 1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [15, 2]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [23, 24, 5]. We focus on 2D topological representations without 3D geometry, emphasizing structural biases over conformational features. We adopt Graphormer-inspired structural biases (shortest-path distance (SPD) + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks as a locality prior for scalability on larger graphs. We treat this mask primarily as a locality prior and efficiency knob; the main methodological contribution is the cardinality-preserving attention (CPA) channel, building on prior CPA work (Zhang & Xie, 2020) and adapted here with query-conditioning and transformer-specific design. The CPA channel retains dynamic support-size signals complementary to static centrality embeddings.
@@
  • A graph transformer layer with query-conditioned cardinality-preserving attention aligned to structured supports, including theoretical motivation connecting to multiset discrimination and expressivity.
  • Dual-objective self-supervised pretraining on $\sim$28M molecules with unified architecture across all ablations.
  • Rigorous matched-protocol evaluation showing consistent, often significant gains over reproduced baselines, expanded to include Graphormer, D-MPNN, GIN, and contextualized against pretrained 2D models (GROVER, MolCLR, GEM). Code and artifacts are provided in the accompanying reproducibility package (reserved DOI: 10.5281/zenodo.18622116; public DOI landing-page access activated upon acceptance if required by venue policy).
@@ 2 Related Work
Message passing remains a dominant paradigm for molecular graphs [4], and attention-based GNNs are widely used [18]. Graph transformers such as Graphormer inject structural biases (shortest-path distance, edge features, centrality) into global attention [23]. GPS [14], a recent hybrid approach, combines local message-passing layers with global attention, serving as a close architectural relative; it demonstrates the value of blending local and global mechanisms. Other transformer variants include Spectral Attention Networks (SAN) [10], which leverage spectral properties for structured attention, and Exphormer [17], which provides a sparse attention framework for graphs. TokenGT [9] proposes tokenized graph representations, while Transformer-M [12] extends transformers to handle both 2D and 3D molecular data, emphasizing multimodal representations. Sparse variants often restrict attention to local neighborhoods (e.g., $K$-hop masking) to reduce computation in practice. Cardinality-preserving attention (CPA) adds unnormalized channels to retain degree/cardinality signals, originally proposed by Zhang & Xie (2020) for MPNNs [26]. Self-supervised pretraining on graphs includes contrastive methods (GraphCL) and masked modeling (GraphMAE) [24, 5]. CardinalGraphFormer unifies structured sparse attention with aligned cardinality preservation and dual-objective pretraining, differing from prior CPA by integrating query-conditioned gating within transformer attention.
Note on 3D Methods. Our work focuses exclusively on 2D topological features without 3D geometry or conformational information. While 3D graph neural networks (e.g., NequIP, Allegro, Equivariant Graph Neural Networks) achieve strong results on geometry-dependent tasks, our structural biases target connectivity patterns and are not designed for 3D data. We position our contribution as complementary rather than competitive with 3D methods, and benchmark against 2D baselines to ensure fair comparison.
Graph Transformer Taxonomy. Graph transformers span three archetypal designs: (1) Local, leveraging sparse message-passing neighborhoods; (2) Global, applying full pairwise attention; (3) Hybrid, combining local aggregation with global interaction (e.g., GPS). Our approach fits the hybrid category, using $K$-hop sparsity for local structure and CPA for dynamic cardinality signals. This taxonomy helps position our contribution relative to other structured attention mechanisms.
@@
3 Theoretical Analysis
We analyze CPA-augmented attention through the lens of multiset expressivity and the Weisfeiler-Leman (WL) framework. Our results formalize how unnormalized aggregation channels recover discriminative power lost by softmax normalization.
3.1 Notation and Setup
Fix a node $i$ with support set $\mathcal{S}(i)\subseteq V$ of cardinality $|\mathcal{S}(i)|=n_{i}$. Let $\mathbf{v}_{j}\in\mathbb{R}^{d}$ denote value vectors, $\mathbf{q}_{i}\in\mathbb{R}^{d}$ the query vector, and $a_{ij}\in\mathbb{R}$ attention logits. The softmax attention output is:
$\mathbf{o}_{i}^{\mathrm{attn}}=\sum_{j\in\mathcal{S}(i)}\alpha_{ij}\mathbf{v}_{j},\quad\alpha_{ij}=\frac{\exp(a_{ij})}{\sum_{k\in\mathcal{S}(i)}\exp(a_{ik})}.$ (1)
The CPA-augmented output is:
$\mathbf{o}_{i}^{\mathrm{CPA}}=\sum_{j\in\mathcal{S}(i)}\alpha_{ij}\mathbf{v}_{j}\;+\;\mathbf{g}_{i}\odot\sum_{j\in\mathcal{S}(i)}\mathbf{v}_{j},\quad\mathbf{g}_{i}=\sigma(W_{g}\mathbf{q}_{i}),$ (2)
where $\sigma(\cdot)$ is the element-wise sigmoid, $W_{g}\in\mathbb{R}^{d\times d}$ is learnable, and $\odot$ denotes the Hadamard product.
3.2 Cardinality Blindness and Its Resolution
Proposition 1 (Cardinality Blindness of Softmax Attention).
Consider two nodes $i$ and $i^{\prime}$ with support sets of different cardinalities $n_{i}\neq n_{i^{\prime}}$, and suppose the following conditions hold:
  1. (C1)
    Identical empirical value distributions: the multiset $\{\!\{\mathbf{v}_{j}:j\in\mathcal{S}(i)\}\!\}$ is a $\lambda$-fold replication of some base multiset $\mathcal{B}$, and $\{\!\{\mathbf{v}_{j^{\prime}}:j^{\prime}\in\mathcal{S}(i^{\prime})\}\!\}$ is a $\lambda^{\prime}$-fold replication of the same $\mathcal{B}$, with $\lambda\neq\lambda^{\prime}$ (hence $n_{i}=\lambda|\mathcal{B}|$, $n_{i^{\prime}}=\lambda^{\prime}|\mathcal{B}|$).
  2. (C2)
    Matching attention profiles: the attention weights respect the replication structure, i.e., for each copy of a value $\mathbf{b}\in\mathcal{B}$, the corresponding $\alpha_{ij}$ sum to the same proportion in both supports. Concretely, $\sum_{j:\mathbf{v}_{j}=\mathbf{b}}\alpha_{ij}=\sum_{j^{\prime}:\mathbf{v}_{j^{\prime}}=\mathbf{b}}\alpha_{i^{\prime}j^{\prime}}$ for every $\mathbf{b}\in\mathcal{B}$.
Then $\mathbf{o}_{i}^{\mathrm{attn}}=\mathbf{o}_{i^{\prime}}^{\mathrm{attn}}$.
Proof.
Under (C2), the softmax output at node $i$ is:
$\mathbf{o}_{i}^{\mathrm{attn}}=\sum_{\mathbf{b}\in\mathcal{B}}\left(\sum_{j:\mathbf{v}_{j}=\mathbf{b}}\alpha_{ij}\right)\mathbf{b}=\sum_{\mathbf{b}\in\mathcal{B}}p_{\mathbf{b}}\cdot\mathbf{b},$
where $p_{\mathbf{b}}=\sum_{j:\mathbf{v}_{j}=\mathbf{b}}\alpha_{ij}$ is the total probability mass assigned to value $\mathbf{b}$. By (C2), $p_{\mathbf{b}}$ is identical for nodes $i$ and $i^{\prime}$. Hence $\mathbf{o}_{i}^{\mathrm{attn}}=\mathbf{o}_{i^{\prime}}^{\mathrm{attn}}$, despite $n_{i}\neq n_{i^{\prime}}$. ∎
Interpretation. Softmax attention computes a weighted average over distinct value identities, discarding how many copies of each value exist. The canonical example is uniform attention: if $\alpha_{ij}=1/n_{i}$ for all $j$, the output is the empirical mean $\bar{\mathbf{v}}$, which is invariant to the number of repetitions. This is the precise sense in which softmax attention is “cardinality-blind.”
Proposition 2 (CPA Resolves Cardinality Blindness).
Under the same conditions (C1)–(C2) of Proposition 1, suppose additionally:
  1. (C3)
    The mean value is nonzero: $\bar{\mathbf{v}}=\frac{1}{|\mathcal{B}|}\sum_{\mathbf{b}\in\mathcal{B}}\mathbf{b}\neq\mathbf{0}$.
  2. (C4)
    The gate has at least one nonzero component: there exists a coordinate $r$ such that $\mathbf{g}_{i}]_{r}>0$.
Then $\mathbf{o}_{i}^{\mathrm{CPA}}\neq\mathbf{o}_{i^{\prime}}^{\mathrm{CPA}}$.
Proof.
The softmax terms are equal by Proposition 1. The unnormalized sums are:
$\sum_{j\in\mathcal{S}(i)}\mathbf{v}_{j}=\lambda\sum_{\mathbf{b}\in\mathcal{B}}\mathbf{b}=\lambda|\mathcal{B}|\bar{\mathbf{v}},\qquad\sum_{j^{\prime}\in\mathcal{S}(i^{\prime})}\mathbf{v}_{j^{\prime}}=\lambda^{\prime}|\mathcal{B}|\bar{\mathbf{v}}.$
Since $\lambda\neq\lambda^{\prime}$, these sums differ. For the gated terms: even if $\mathbf{g}_{i}=\mathbf{g}_{i^{\prime}}$ (which holds when the queries are identical), we have at coordinate $r$:
$\mathbf{g}_{i}]_{r}\cdot\lambda|\mathcal{B}|[\bar{\mathbf{v}}]_{r}\neq[\mathbf{g}_{i^{\prime}}]_{r}\cdot\lambda^{\prime}|\mathcal{B}|[\bar{\mathbf{v}}]_{r}$
since $\mathbf{g}_{i}]_{r}>0$, $\bar{\mathbf{v}}]_{r}\neq 0$ (some coordinate must be nonzero by (C3)), and $\lambda\neq\lambda^{\prime}$. Hence $\mathbf{o}_{i}^{\mathrm{CPA}}\neq\mathbf{o}_{i^{\prime}}^{\mathrm{CPA}}$. ∎
3.3 Connection to the 1-WL Framework
We connect CPA to the Weisfeiler-Leman (1-WL) color refinement algorithm, building on the characterization by Xu et al. [21] and Morris et al. [13].
Theorem 1 (CPA-Augmented Attention and 1-WL Expressivity).
Let $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{D}$ be an injective function (realizable, e.g., by a sufficiently wide MLP with generic weights). Consider the composed CPA aggregation:
$h_{i}=\psi\!\left((1+\epsilon)\,\varphi(\mathbf{x}_{i})+\mathbf{g}_{i}\odot\sum_{j\in\mathcal{S}(i)}\varphi(\mathbf{x}_{j})\right),$ (3)
where $\mathbf{x}_{j}$ are node features, $\epsilon>0$ is a learnable scalar, and $\psi:\mathbb{R}^{D}\to\mathbb{R}^{D^{\prime}}$ is an injective function (another MLP). Suppose the gate satisfies $\mathbf{g}_{i}\succ\mathbf{0}$ component-wise (achievable for generic $W_{g}$ since sigmoid is strictly positive). Then $h_{i}$ is an injective function of the pair $(\mathbf{x}_{i},\{\!\{\mathbf{x}_{j}:j\in\mathcal{S}(i)\}\!\})$, and consequently one layer of this aggregation is at least as expressive as one iteration of 1-WL color refinement.
Proof.
By the characterization in Xu et al. [21] (Theorem 3 and Lemma 5), an aggregation scheme achieves 1-WL expressivity if and only if the aggregation function over the neighbor multiset is injective and composed with the self-feature via an injective function. The critical requirement is that the multiset aggregation be injective.
Step 1: Injectivity of gated sum aggregation. Since $\mathbf{g}_{i}\succ\mathbf{0}$, the map $\mathbf{z}\mapsto\mathbf{g}_{i}\odot\mathbf{z}$ is a bijection on $\mathbb{R}^{D}$ (with inverse $\mathbf{z}\mapsto\mathbf{g}_{i}^{-1}\odot\mathbf{z}$, where inversion is element-wise). Therefore, $\mathbf{g}_{i}\odot\sum_{j}\varphi(\mathbf{x}_{j})$ is injective as a function of the multiset $\{\!\{\varphi(\mathbf{x}_{j})\}\!\}$ if and only if $\sum_{j}\varphi(\mathbf{x}_{j})$ is.
Step 2: Injectivity of sum over embedded features. By Xu et al. [21] (Lemma 5, building on results of Zaheer et al.), for any countable multiset domain $\mathcal{X}\subset\mathbb{R}^{d}$, there exists a mapping $\varphi$ and a function $f$ such that $f\!\left(\sum_{\mathbf{x}\in\mathcal{M}}\varphi(\mathbf{x})\right)$ is unique for each multiset $\mathcal{M}$. A sufficiently wide MLP with generic weights realizes such a $\varphi$ over bounded feature domains.
Step 3: Composing with self-features. The term $(1+\epsilon)\,\varphi(\mathbf{x}_{i})$ injects the center node’s features with a distinct scale ($\epsilon>0$ prevents cancellation with the sum term). The outer function $\psi$ preserves injectivity by assumption. Together, $h_{i}$ is injective in $(\mathbf{x}_{i},\{\!\{\mathbf{x}_{j}\}\!\})$, matching one WL iteration.
Connection to CPA in practice. In CardinalGraphFormer, the value projection $V:\mathbf{x}_{j}\mapsto\mathbf{v}_{j}$ plays the role of $\varphi$, and the feed-forward network following attention plays the role of $\psi$. The CPA output contains the sum $\sum_{j}\mathbf{v}_{j}$ as an additive channel, providing the necessary injective multiset aggregation alongside the softmax channel. While the actual learned weights may not exactly satisfy the generic-weight conditions, the architecture has the capacity for 1-WL expressivity, unlike pure softmax attention which structurally lacks it. ∎
Corollary 1 (Degree-Normalized CPA Has Reduced Expressivity).
Consider the degree-normalized variant:
$\mathbf{o}_{i}^{\mathrm{Norm}}=\sum_{j\in\mathcal{S}(i)}\alpha_{ij}\mathbf{v}_{j}+\mathbf{g}_{i}\odot\frac{1}{|\mathcal{S}(i)|}\sum_{j\in\mathcal{S}(i)}\mathbf{v}_{j}.$ (4)
The normalized aggregation $\frac{1}{|\mathcal{S}(i)|}\sum_{j}\mathbf{v}_{j}$ is not injective over multisets: there exist distinct multisets $\mathcal{M}\neq\mathcal{M}^{\prime}$ with identical means. Consequently, Norm-CPA cannot achieve 1-WL expressivity in general.
Proof.
Mean aggregation maps different multisets to the same value whenever the multisets share the same centroid. A concrete counterexample: let $\mathcal{M}=\{\!\{\mathbf{v}\}\!\}$ (a single element) and $\mathcal{M}^{\prime}=\{\!\{\mathbf{v},\mathbf{v}\}\!\}$ (two copies). Then $\frac{1}{1}\mathbf{v}=\frac{1}{2}(2\mathbf{v})=\mathbf{v}$. More generally, for any multiset $\mathcal{M}$ with mean $\bar{\mathbf{v}}$, the multiset $\mathcal{M}\cup\{\!\{\bar{\mathbf{v}}\}\!\}$ has the same mean. Since injectivity of the neighbor aggregation is necessary for 1-WL (Xu et al. [21], Theorem 3), Norm-CPA cannot generally match 1-WL. ∎
Remark 1 (Scope and Limitations).
We state the scope of the above results for transparency.
  1. 1.
    Per-layer analysis. All results characterize single-layer expressivity. Multi-layer CPA composes injective functions (which preserves injectivity), but shared parameters across nodes and the cumulative effect of sigmoid bounding may introduce practical constraints not captured here.
  2. 2.
    Sigmoid bounding. The gate $\mathbf{g}_{i}\in(0,1)^{d}$ is strictly positive (sigmoid never reaches zero), satisfying the technical requirement $\mathbf{g}_{i}\succ\mathbf{0}$. However, values near zero attenuate the unnormalized channel, potentially weakening expressivity in practice. We note that unbounded gates caused training instability in our experiments (Appendix, Table 15).
  3. 3.
    Existence vs. learnability. Theorem 1 is an existence result: appropriate weight configurations exist, but gradient-based optimization may not find them. Our empirical results (Table 13) provide evidence that useful configurations are reachable in practice, though the gap between theoretical capacity and learned behavior remains open.
  4. 4.
    Countable domain assumption. The injectivity result for sum aggregation (Xu et al. [21], Lemma 5) formally holds over countable domains. Molecular graphs, with discrete atom and bond features, satisfy this condition for the input layer; subsequent layers operate in continuous space where the result holds generically (with probability 1 over random weight initialization).
  5. 5.
    Relation to higher-order WL. CPA enhances 1-WL expressivity but does not address higher-order ($k$-WL) limitations. Graphs indistinguishable by 1-WL (e.g., certain regular graphs) remain indistinguishable by CPA-augmented attention. Higher-order positional encodings or subgraph-based methods would be needed for such cases.
@@ 4.2 Structured Sparse Attention with Graphormer-Inspired Biases
The support set is $\mathcal{S}(i)=\{j\mid\mathrm{SPD}(i,j)\leq K\}$ (includes self, SPD(i,i)=0(i,i)=0 meaning the self-attention position uses the $\phi(\mathrm{SPD}(i,i))=0$ bias bin explicitly). Main experiments use $K=3$. Attention is masked to $\mathcal{S}(i)$ (outside: -$\infty$). Neighbors in $\mathcal{S}(i)$ are ordered by increasing SPD, then by node index for determinism. Logits:
@@
We implement structured attention by precomputing neighbor index lists per node and using gather-attention-scatter operations over the support set. To remain compatible with optimized attention kernels (e.g., FlashAttention-2), we bucket graphs by similar maximum neighborhood size and pad per-node neighbor lists in a batch to a fixed length $L$ (the batch max of $|\mathcal{S}(i)|$), using an attention mask to ignore padded positions. Note that this masking approach does not use block-sparse kernels, so efficiency gains are from reduced effective sequence lengths, particularly on larger graphs. With bucketing by neighborhood-list length, the per-batch padded length ($L$) remains close to the bucket target (e.g., mean $L\approx 25$ on the full corpus), yielding $10\%$ average padding.
@@ 4.3 Cardinality-Preserving Attention
[…]
Figure 1: Compact comparison of a standard attention head and the CPA-augmented head. CPA keeps the same support set $\mathcal{S}(i)$ and normalized path, then adds a query-gated unnormalized support-sum branch before output projection.
Figure 1 summarizes this parallel-path formulation and highlights that CPA changes aggregation rather than support definition.
The degree-normalized CPA ablation (Table 13) divides the unnormalized sum by $|\mathcal{S}(i)|$ before gating and underperforms, consistent with the hypothesis that raw magnitude variation—correlated with current support size—carries useful information beyond static degree bins or explicit normalization. Supporting this, the L2 norm of the gated unnormalized sum (averaged across heads and layers, post-LayerNorm) correlates positively with $|\mathcal{S}(i)|$ even after controlling for node degree (partial correlation $r=0.52\pm 0.03$ (95% CI: [0.46, 0.58])), computed on a held-out 50k-graph pretraining validation subset, averaged across 5 seeds.
@@ 4.4 Self-Supervised Objectives
Contrastive views. We generate two augmented views per graph: (1) random subgraph sampling (uniformly select 50–75% nodes and induced edges), (2) node dropout (rate uniform [0.1,0.3] on surviving nodes), (3) edge dropout (rate uniform [0.1,0.3] on surviving edges). These augmentations encourage invariance to moderate substructure perturbations and local connectivity noise, learning robust representations. To address conceptual mismatch concerns, we recompute SPD distances and degree bins per view as the PRIMARY augmentation strategy, intersecting $\mathcal{S}(i)$ with surviving nodes. This increases computation but enhances semantic consistency. We also evaluate chemistry-aware alternatives (attribute masking, valency-constrained edge perturbations) as ablations (Table 1), showing CPA gains persist with both random and chemistry-aware augmentations.